OpenStudy (errinss):

Alice purchased paint in a bucket with a radius of 9 in. and a height of 9.5 in. The paint cost $0.09 per in3. What was the total cost of the paint, rounded to the nearest cent? Use 3.14 to approximate pi.

OpenStudy (hauntedwoodsgal):

V=π×r2×h where r is the radius of our buck, and h is the heoght of our buck. so what is the volume V?

OpenStudy (errinss):

2,416.23?

OpenStudy (hauntedwoodsgal):

Yes!!!

OpenStudy (errinss):

woohoo!!

OpenStudy (wolf1728):

Corrrection: Cone Volume = (π • r² • h) ÷ 3

OpenStudy (errinss):

this is a cylinder....

OpenStudy (wolf1728):

Sorry :-(

OpenStudy (errinss):

nah your good

OpenStudy (wolf1728):

thanks :-)

OpenStudy (errinss):

Kirsten used a log to make an umbrella stand. The log was 28 in. tall and had a radius of 7.5 in. Kirsten cut a cylindrical shape, with a radius of 5.5 in., out of the center of the log. After Kirsten cuts out the cylinder, what is the volume of the remaining portion of the log? Use 3.14 to approximate pi, and express your answer to two decimal places.

OpenStudy (hauntedwoodsgal):

First find the volume of the whole log.

OpenStudy (hauntedwoodsgal):

V=πr2h

OpenStudy (wolf1728):

Volume = PI * r^2 * height Volume = PI * 5.5^2*28

OpenStudy (hauntedwoodsgal):

Plug in what we know: \(V = (3.14)(7.5^2)(28)\ Multiply 3.14 * 7.5 * 7.5 * 28

OpenStudy (errinss):

4,945.5

OpenStudy (hauntedwoodsgal):

yeasss

OpenStudy (wolf1728):

I think they want the volume of the remaining portion (radius = 5.5)

OpenStudy (hauntedwoodsgal):

Oh?

OpenStudy (hauntedwoodsgal):

Multiply 3.14 * 5.5 * 5.5 * 28

OpenStudy (errinss):

2,659.58

OpenStudy (errinss):

Joan’s washing machine tub is a cylinder with a height of 17 in. and a diameter of 22 in. Her washer was half full of water when it stopped working. The repair person’s pump removed 46.2 in3 of water every second. How many seconds did it take to empty the washer? Use 3.14 to approximate pi, and round your answer to the nearest whole number. A. 70 s B. 140 s C. 280 s D. 559 s

OpenStudy (hauntedwoodsgal):

And what do you think it is?

OpenStudy (errinss):

B....

OpenStudy (wolf1728):

radius = 11 inches Volume = PI * 11*11* (17/2) tub was half filled with water

OpenStudy (errinss):

a?

OpenStudy (wolf1728):

I think "a" is the correct answer

OpenStudy (hauntedwoodsgal):

I agree @wolf1728

OpenStudy (errinss):

A large cylinder contains a saline solution for science experiments. The cylinder has a height of 40 mm and a radius of 12 mm. Braddock pours the saline solution into test tubes that each hold 3014 mm3. How many test tubes can Braddock prepare? Use 3.14 to approximate pi, and round your final answer to the nearest whole number.

OpenStudy (wolf1728):

Do you have an answer yet errinss?

OpenStudy (errinss):

uhm..i think its 6

OpenStudy (wolf1728):

That's what I have - we must be right Very good!

OpenStudy (errinss):

woohoo, one more?

OpenStudy (wolf1728):

Okay

OpenStudy (errinss):

Monte performs an experiment using 2 identical graduated cylinders with a radius of 2 cm. The volume of the liquid in the first graduated cylinder is 188.4 cm3. The volume of the liquid in the second graduated cylinder is 314 cm3. What is the difference in the height of the liquid in the two cylinders? Use 3.14 to approximate pi.

OpenStudy (errinss):

i think 10cm is the difference

OpenStudy (errinss):

Question 1 Incorrect Question 2 Incorrect Question 3 Correct Question 4 Correct Question 5 Correct

OpenStudy (wolf1728):

Since V= PI*r^2*height 1st cyl 188.4 = PI * 2*2 * height height = 15 2nd cyl 314 = PI * 2*2* height height = 25 Difference = 10 cm Very good errinss!!!

OpenStudy (errinss):

i got 60% :(

OpenStudy (wolf1728):

Did you use all the answers you posted here?

OpenStudy (errinss):

ya

OpenStudy (wolf1728):

geez, I thought you had done those correctly.

OpenStudy (errinss):

1. 217.46 2. 2285.92

OpenStudy (wolf1728):

For the first one I got 217.57 (I'm using PI= 3.14159265...)

OpenStudy (errinss):

ya thats what your supposed to use i dont know what i did but i got it wrong

OpenStudy (errinss):

The East-Side soccer team has a cylindrical water cooler that is half full. The diameter of the cooler is 12 in. and the height is 15 in. The team’s drinking cups hold 20.14 in3. How many drinking cups can be filled from the cooler? Use 3.14 to approximate pi, and round your answer to the nearest whole number. A. 14 B. 42 C. 84 D. 169

OpenStudy (errinss):

would it be C?

OpenStudy (wolf1728):

No - don't forget it is half-full

OpenStudy (errinss):

A

OpenStudy (wolf1728):

No it would be "B"

OpenStudy (errinss):

oh...

OpenStudy (errinss):

A bricklayer uses a cylinder to store sand. The cylinder is 5 ft tall and has a radius of 0.5 ft. Each batch of mortar contains 0.48 ft3 of sand. How many batches of mortar can the bricklayer make with 1 cylinder of sand? Use 3.14 to approximate pi, andround your final answer to the nearest whole number.

OpenStudy (wolf1728):

Cylinder volume = PI*.5*.5*5 = 3.927 cubic feet

OpenStudy (errinss):

haha thats waht i got...kinda i got 3.925

OpenStudy (wolf1728):

3.927 / .48 = 8.812 Round that to 8 batches

OpenStudy (errinss):

ok there is one more question

OpenStudy (wolf1728):

Well let's get to that one

OpenStudy (errinss):

A dispenser holds 5652 cm3 of liquid soap and is now full. The radius of the dispenser is 7.5 cm. What is the difference between the height of the soap in the full dispenser and the height when 4239 cm3 of soap remains in the dispenser? Use 3.14 to approximate pi. im thinking its 30 cm is the difference

OpenStudy (wolf1728):

Height = volume / PI*r*r Height = 5,652 / 176.71 Height = 31.98 = 32 (rounded) (when full)

OpenStudy (errinss):

i still got 60% wth ugh

OpenStudy (errinss):

#2 and #5

OpenStudy (errinss):

Question 1 Correct Question 2 Incorrect Question 3 Correct Question 4 Correct Question 5 Incorrect

OpenStudy (wolf1728):

Height = 4,239 / 176.71 Height = 23.98 = 24 (rounded) Difference = 8

OpenStudy (wolf1728):

Is question 5 the one we just did?

OpenStudy (errinss):

YA

OpenStudy (errinss):

i accidentally hit 9 my finger slipped

OpenStudy (wolf1728):

oh geez one slip makes it wrong!

OpenStudy (wolf1728):

And what was question #2?

OpenStudy (errinss):

lol uhm i cant remember but i did that one without you lol :( my fault

OpenStudy (wolf1728):

sheesh well maybe it's time to end this - for now

OpenStudy (errinss):

for now...lol okay

OpenStudy (wolf1728):

Okay - cya

OpenStudy (errinss):

bye thank you for your help

OpenStudy (wolf1728):

u r welcome